Alvin is 3 years older than his younger sister, Betty. When Alvin was half the age he is now, Betty was 9 yrs older than half his age at that time. How old are Alvin and Betty?
Answer by hamsanash1981@gmail.com(151)

Let Betty's age be x and Alvins age = x+3
When Alvin was half the age , Betty was 9 yrs older than half his age at that time then
1/2(x+3) = x-9
x+3 = 2x-18
x = 21
Alvins age = 24 and betty's age is 21